# Language Models, Graph Searching, and Supervision Adulteration: When More Supervision is Less and How to Make More More

### v1

@@ B.2 Learnability of Graphs
B.3 Spurious Correlations and Shortcuts
@@
Wang et al. (2023b) showed that spurious correlations affect the performance of LLMs on graph tasks. In particular, they design two special types of graphs; a ‘chain’ which is just a very long path and a ‘clique’ which has a high edge density. They found that LLMs fail to solve a connectivity task at high rates on these graphs compared to other general graphs. This implies that the underlying algorithm is not learnt (or being applied consistently across different graph types) and thus a shortcut is being employed. (Jiang et al., 2024a; Mirzadeh et al., 2025) showed similar results for reasoning tasks where they argue that the model is learning in-domain spurious correlations and thus only learning a superficial pattern matching instead of true reasoning. Press et al. (2023) showed that LLMs can often correctly solve multi-hop subtasks without getting the overall or final answer correct, which they attribute to fact memorization which can be considered as a spurious correlation or undesired shortcut.
@@
B.4 Masking
Masking is often done to avoid spurious correlations and overfitting. Masks can be crafted or structured depending on the task via inductive biases that mask specifically linked tokens. Deng et al. (2021) used constructed query-evidence data pairs and a masked spanning objective that masks parts of the query that are supported by evidence, thus inducing the model to learn a connection between the evidence and the query. Span selection is also needed in other ways of supervised training of reasoning tasks (Stacey et al., 2022). Rabe et al. (2021) used masking for math reasoning by masking specific sub-expressions. This used an inductive bias which masks all occurrences of such sub-expression. Chen et al. (2024a) showed that masking tokens within the CoT improved their effectiveness for fine-tuning and that the placement of the masking is important.
@@ B.4 Masking
B.5 Expressivity and Learnability
Various works have considered the computational limits or expressivity of transformers i.e. ‘can a transformer actually solve this problem’ (given a particular capacity in terms of hidden-state size or number of layers) (Yun et al., 2020). Various computational models are used to prove expressibility, such as formal logic (Merrill and Sabharwal, 2023), formal languages (Hao et al., 2022; Strobl et al., 2024b), massively parallel computation (Sanford et al., 2024b), or declarative programming languages such as RASP (Restricted Access Sequence Programming)(Weiss et al., 2021).
Weiss et al. (2021) demonstrated that RASP programs upper-bound the difficulty/complexity of a task (for a transformer) in terms of the number of required layers (and attention heads) required to solve the task. It employs a limited computational model of transformers that are restricted to performing uniform attention over a subset of queries (average-hard attention (Strobl et al., 2024b)). [^21] See Yang et al. (2024) who consider when soft attention can simulate various kinds of hard attention. While this excludes RASP’s use to model numerical tasks, it does make it easy to model symbolic tasks such as path-star. Zhou et al. (2024a) extended RASP to causal attention and conjecture that short RASP programs lead to length-generalizability. [^22] They also wrote RASP in Numpy, making it an easy tool for NLP/ML practitioners. Huang et al. (2025) formalized this conjecture, showing why certain problems have poor length generalization while also showing that a certain class of tasks have guaranteed length generalization. Strobl et al. (2024a) extended RASP to model transformers as transducers, which requires accounting for non-length preserving transitions. RASP programs can be compiled into actual transformers and the reverse (Friedman et al., 2023; Lindner et al., 2023).
@@ B.5 Expressivity and Learnability
de Luca and Fountoulakis (2024) showed that looped transformers can express various graph algorithms with a constant number of layers. They used a modified transformer architecture which allows for encoding a graph as an adjacency matrix, with a special attention mechanism over this matrix. They made significant note of the need to limit numerical errors through various methods like using hard attention and careful choice of positional embeddings (see Appx. B.7). Sanford et al. (2024a) developed a representational hierarchy of problem classes for transformers on graph problems. Path-star falls under the ‘parallelizable tasks’ class, in particular, those solvable with logarithmic depth. Frydenlund (2024) showed that transformers can express the path-star task via RASP for encoder- and decoder-only models.
@@
Expressibility is not to be confused with learnability, i.e. ‘can standard learning methods be used to train a transformer to solve this problem’ (Allen-Zhu and Li, 2023; Deletang et al., 2023; Sanford et al., 2023, 2024b). [^23] See Svete and Cotterell (2024) and Svete et al. (2024) for a case study, where the former considered the expressibility of transformers to model $N$-gram language models, and develop various computational models either using $N-1$ layers or $N-1$ attention heads in combination with hard/sparse attention, while the later then considered the learnability of such models. Going back to RASP, Zhou et al. (2024a) modified RASP to better model numerical representation and align RASP with empirical results about learnability. This included only allowing single increment indexing. Chang and Bisk (2025) pointed out that transformers fail to count inductively and as such, such abilities should not be inherent abilities in the computational model. This was also an argument stemming from empirical learnability results. This demonstrates an inherent divide between the models used for expressibility and learnability.
@@
Learnability is the core question of this work i.e. can decoder-only transformers learn the path-star task? We show this empirically as well as provide theoretical explanations for why adulteration or lack of decomposition causes the task to be unlearnable.
@@
de Luca and Fountoulakis (2024) also considered a small number of learnability experiments using the CLRS dataset. Here they train on 16 node graphs and evaluate on 64 node graphs as a form of length generalization (see Appx. B.7). They highlighted how learnability is much more difficult than expressibility where ‘despite demonstrating the existence of parameters capable of [graph] simulation, discovering them through gradient-based training is challenging.’
B.6 Sensitivity
@@
A specific and highly relevant case of transformer expressivity and learnability is for parity due to being a (maximally) sensitive function (Hahn et al., 2021; Bhattamishra et al., 2023; Hahn and Rofin, 2024). [^24] Again, see Hu et al. (2025a), for a connection between parity and the path-star task. The sensitivity of a discrete function on an input sequence $x$ describes the number of disjoint subsets of $x$ which, when changed, cause changes to the output. Thus functions with low sensitivity contain redundant information across $x$, whereas functions with high sensitivity have tokens that isolate important information. The path-star task completely changes its output based on a single target token provided in the query. Another view of sensitivity is as an analog to the smoothness of continuous functions, where path-star is not smooth with respect to a change in target.
@@
B.7 Length Generalization, Task Decomposition, and Scratchpads
The effect of task decomposition on learnability has been studied (Wies et al., 2023; Dziri et al., 2023; Abbe et al., 2024b). This is often studied in the context of length generalization. This is a specific kind of OOD generalization of great importance to LMs due to their sequential nature (Anil et al., 2022; Zhou et al., 2024a, b). Length generalization relates to reasoning tasks that scale to the number of required reasoning steps (hops) (Dziri et al., 2023; Abbe et al., 2024a; Xiao and Liu, 2024, 2025; Mirzadeh et al., 2025). One of the main concerns about the difficulty for transformers to learn parity is that, when they do learn the task, this does not generalize to unseen sequence lengths (Bhattamishra et al., 2020; Hahn and Rofin, 2024). This betrays the fact that the underlying algorithm has not been learnt by the model, i.e. “the failure in length generalization corroborates the models’ fundamental limitation that they may not genuinely understand the task solving algorithm but may rely on short-cut learning that is only applicable to sequences of trained length” (Cho et al., 2024b). In addition to parity, integer addition is often used as a test-bed for length generalization (McLeish et al., 2024; Cho et al., 2024b). Chang and Bisk (2025) showed that transformers show poor OOD performance for the simple task of counting (including task variants). Deletang et al. (2023) showed similar results for a series of more challenging tasks based on formal languages grouped within the Chomsky hierarchy.
@@ B.7 Length Generalization, Task Decomposition, and Scratchpads
Like the path-star task, integer arithmetic is a simple task with a simple underlying algorithm that LMs fail on. Also, a trivial reverse solution exists. Zhou et al. (2024b) shows that the reverse solution to arithmetic is more robust in terms of length generalization. Note that both these reverse solutions work because they do not need to think multiple steps ahead. For the path-star task, this betrays a lack of reasoning, however, for addition, this conforms with how humans perform arithmetic and thus feels like a more valid solution despite not requiring multi-step reasoning. The more interesting question is if models can learn to find the trivial order (which we find does not happen for the BoW experiments). See Appx. B.9 for order considerations.
@@
Yin et al. (2024) tried thinking tokens for the path-star problem to negative results. [^27] Note these are done under the original settings that allow for spurious correlations, and thus may have failed due to other reasons. This is interesting because using $M$ thinking tokens can, in theory, provide a trivial solution by computing the reverse arm with the thinking token and then the forward arm from the reverse solution (similar to the BoW experiments, just with even less supervision). We conjecture that thinking tokens do not work for the path-star task as they do not provide additional decomposition supervision.
B.8 Positional Embeddings
With length generalization comes the need to have positional embeddings that allow for exact matching across long lengths and generalize to unseen positions (Kiyono et al., 2021; Kazemnejad et al., 2023; Ruoss et al., 2023; Li et al., 2024a; McLeish et al., 2024). Chang and Bisk (2025) showed that different embedding types generalize differently to different counting tasks. Such methods will be important considerations for graph-based tasks when scaling up the size of graphs and considering length generalization. This can also lead to some unexpected results like using no positional embeddings (NoPe) being possible for decoder-only models (Irie et al., 2019; Tsai et al., 2019; Haviv et al., 2022; Chi et al., 2023; Kazemnejad et al., 2023; Wang et al., 2024b; Irie, 2024; Zuo et al., 2025) and can lead to better length generalization for symbolic reasoning tasks (Kazemnejad et al., 2023). However, Wang et al. (2024b) showed that NoPe fails to generalize due to a collapse in the attention head distribution as the context size increases. (Yang et al., 2025) followed this up with a hybrid strategy that combines NoPe, for its strong token retrieval and RoPe for its inductive biases. Frydenlund (2024) found that the choice of positional embedding mattered for the path-star task and that NoPe worked when using decoder-only models.
@@ B.8 Positional Embeddings
B.9 Order Matters and Reversal Curse
Prior work has explored the impact of ordering source-side information in LLMs for premise order on the complex tasks of reasoning (Wang et al., 2023a; Chen et al., 2024c; Allen-Zhu and Li, 2024; Shah et al., 2024) and proof generation (An et al., 2024). Liu et al. (2024) has shown that LLMs struggle to retrieve relevant information in long contexts when that information is placed in the middle of the context compared to either the front or back. Frydenlund (2024) showed that order matters for the path-star where they considered that the query should proceed the graph.
@@ B.9 Order Matters and Reversal Curse
Papadopoulos et al. (2024) discovered an asymmetry in perplexity between models trained either in the forward or reverse direction, with the forward having consistently lower perplexity. This is surprising given that both directions give a valid and theoretically equivalent decomposition of the sequential probability.
@@
B.10 Non-AR, Iterative-AR, and Discrete Diffusion Models
@@
Bachmann and Nagarajan (2024) used a ‘teacher-less’ model which masks out all input tokens. Frydenlund (2024) connected this model to NAR models and also showed that the path-star task was solvable via an encoder-only model with NAR and IAR training. This was based on a modified version of the CMLM model (where the conditional ‘C’ part of the model is removed) (Ghazvininejad et al., 2019). Frydenlund (2024) incorrectly implied that the original ‘teacher-less’ model was non-causal when considering it as a NAR model. The model described by Monea et al. (2023) is meant to modify an autoregressive model post-hoc and thus is designed to keep the causal constant. However, in terms of independently modeling and predicting multiple tokens (i.e. loss, training, and inference procedure) it behaves exactly like a NAR model.
@@
B.11 Future Token Prediction
Early works in future prediction designed models which could predict $N$ tokens into the future by creating $N$ separate hidden-states and training on each state with cross-entropy against a single future token (Goodman et al., 2020; Qi et al., 2020). Thus these are not truly belief-states directly, however, a belief-state must be present in the model in order to generate the $N$ separate hidden-states. The general goal of this was for improved training by explicitly learning to predict future tokens and hence plan for future tokens, and was not for multi-token inference. Heo et al. (2024) also used multi-state prediction for future n-grams but also introduced a method to explicitly create representations that are compositional into the future. Gloeckle et al. (2024) proposed an efficient training method for $N$ token prediction. Cai et al. (2024) repurposed $N$ multi-head attention to create the $N$ hidden-states and then used this multi-token prediction for faster inference via speculative decoding (Xia et al., 2022; Chen et al., 2023a; Leviathan et al., 2023). [^28] As an aside, speculative decoding was also the original purpose of the ’teacher-less’ model (Monea et al., 2023) and, as with NAR/IAR, the main motivating factor for speculative decoding is improved inference speed..

### v3

@@ C.2 Learnability of Graphs
C.3 Spurious Correlations and Shortcuts
@@
Wang et al. (2023b) showed that spurious correlations affect the performance of LLMs on graph tasks. In particular, they design two special types of graphs; a ‘chain’ which is just a very long path and a ‘clique’ which has a high edge density. They found that LLMs fail to solve a connectivity task at high rates on these graphs compared to other general graphs. This implies that the underlying algorithm is not learnt (or being applied consistently across different graph types) and thus a shortcut is being employed. Jiang et al. (2024a); Mirzadeh et al. (2025) showed similar results for reasoning tasks where they argue that the model is learning in-domain spurious correlations and thus only learning a superficial pattern matching instead of true reasoning. Press et al. (2023) showed that LLMs can often correctly solve multi-hop subtasks without getting the overall or final answer correct, which they attribute to fact memorization which can be considered as a spurious correlation or undesired shortcut.
@@
C.4 Masking
Masking is often done to avoid spurious correlations and overfitting. Masks can be crafted or structured depending on the task via inductive biases that mask specifically linked tokens. Deng et al. (2021) used constructed query-evidence data pairs and a masked spanning objective that masks parts of the query that are supported by evidence, thus inducing the model to learn a connection between the evidence and the query. Span selection is also needed in other ways of supervised training of reasoning tasks (Stacey et al., 2022). Rabe et al. (2021) used masking for math reasoning by masking specific sub-expressions. This used an inductive bias which masks all occurrences of such sub-expression (thus the mask is structured with the task). Chen et al. (2024a) showed that masking tokens within the CoT improved their effectiveness for fine-tuning and that the placement of the masking is important.
@@ C.4 Masking
C.5 Expressivity and Learnability
Various works have considered the computational limits or expressivity of transformers, i.e., ‘can a transformer actually solve this problem’ (given a particular capacity in terms of hidden-state size or number of layers etc.) (Yun et al., 2020). Various computational models are used to prove expressibility, such as formal logic (Merrill and Sabharwal, 2023), formal languages (Hao et al., 2022; Strobl et al., 2024b), massively parallel computation (Sanford et al., 2024b), or declarative programming languages such as RASP (Restricted Access Sequence Programming)(Weiss et al., 2021).
Weiss et al. (2021) demonstrated that RASP programs upper-bound the difficulty/complexity of a task (for a transformer) in terms of the number of required layers (and attention heads) required to solve the task. It employs a limited computational model of transformers that are restricted to performing uniform attention over a subset of queries (i.e., average-hard attention (Strobl et al., 2024b)). [^30] See Yang et al. (2024) who consider when soft attention can simulate various kinds of hard attention. While this excludes RASP’s use on numerical tasks, it does make it easy to model symbolic tasks such as path-star. Zhou et al. (2024a) extended RASP to causal attention and conjecture that short RASP programs lead to length-generalizability. [^31] They also wrote RASP in Numpy, making it an easy tool for NLP/ML practitioners. Huang et al. (2025) formalized this conjecture, showing why certain problems have poor length generalization while also showing that a certain class of tasks have guaranteed length generalization. Strobl et al. (2024a) extended RASP to model transformers as transducers, which requires accounting for non-length preserving transitions. RASP programs can be compiled into actual transformers and the reverse (Friedman et al., 2023; Lindner et al., 2023).
@@ C.5 Expressivity and Learnability
de Luca and Fountoulakis (2024) showed that looped transformers can express various graph algorithms with a constant number of layers. They used a modified transformer architecture which allows for encoding a graph as an adjacency matrix, with a special attention mechanism over this matrix. They made significant note of the need to limit numerical errors through various methods like using hard attention and careful choice of positional embeddings (see Appx. C.7). Sanford et al. (2024a) developed a representational hierarchy of problem classes for transformers on graph problems. Path-star falls under the ‘parallelizable tasks’ class, in particular, those solvable with logarithmic depth. Frydenlund (2024) showed that transformers can express the path-star task via RASP for encoder- and decoder-only models.
@@
Expressibility is not to be confused with learnability, i.e., ‘can standard learning methods be used to train a transformer to solve this problem’ (Allen-Zhu and Li, 2023; Deletang et al., 2023; Sanford et al., 2023, 2024b). [^32] See Svete and Cotterell (2024) and Svete et al. (2024) for a case study, where the former considered the expressibility of transformers to model $N$-gram language models, and develop various computational models either using $N-1$ layers or $N-1$ attention heads in combination with hard/sparse attention, while the later then considered the learnability of such models.
Going back to RASP, Zhou et al. (2024a) modified RASP to better model numerical representation and align RASP with empirical results about learnability. This included only allowing single increment indexing. Chang and Bisk (2025) pointed out that transformers fail to count inductively and as such, such abilities should not be inherent abilities in the computational model. This was also an argument stemming from empirical learnability results. This demonstrates an inherent divide between the models used for expressibility and learnability.
@@
Learnability is the core question of this work, i.e., can decoder-only transformers learn the path-star task? We show this empirically as well as provide theoretical explanations for why adulteration or lack of decomposition causes the task to be unlearnable.
@@
de Luca and Fountoulakis (2024) also considered a small number of learnability experiments using the CLRS dataset. Here they train on 16 node graphs and evaluate on 64 node graphs as a form of length generalization (see Appx. C.7). They highlighted how learnability is much more difficult than expressibility where ‘despite demonstrating the existence of parameters capable of [graph] simulation, discovering them through gradient-based training is challenging.’
C.6 Sensitivity
@@
A specific and highly relevant case of transformer expressivity and learnability is for parity due to it being a (maximally) sensitive function (Hahn et al., 2021; Bhattamishra et al., 2023; Hahn and Rofin, 2024). [^33] Again, see Hu et al. (2025a), for a connection between parity and the path-star task. The sensitivity of a discrete function on an input sequence $x$ describes the number of disjoint subsets of $x$ which, when changed, cause changes to the output. Thus functions with low sensitivity contain redundant information across $x$, whereas functions with high sensitivity have tokens that isolate important information. The path-star task completely changes its output based on a single target token provided in the query. Another view of sensitivity is as an analog to the smoothness of continuous functions, where path-star is not smooth with respect to a change in target.
@@
C.7 Length Generalization, Task Decomposition, and Scratchpads
The effect of task decomposition on learnability has been studied (Wies et al., 2023; Dziri et al., 2023; Abbe et al., 2024b). This is often studied in the context of length generalization. This is a specific kind of OOD generalization of great importance to LMs due to their sequential nature (Anil et al., 2022; Zhou et al., 2024a, b). Length generalization relates to reasoning tasks that scale to the number of required reasoning steps (hops) (Dziri et al., 2023; Abbe et al., 2024a; Xiao and Liu, 2024, 2025; Mirzadeh et al., 2025). One of the main concerns about the difficulty for transformers to learn parity is that, when they do learn the task, this does not generalize to unseen sequence lengths (Bhattamishra et al., 2020; Hahn and Rofin, 2024). This betrays the fact that the underlying algorithm has not been learnt by the model, i.e., “the failure in length generalization corroborates the models’ fundamental limitation that they may not genuinely understand the task solving algorithm but may rely on short-cut learning that is only applicable to sequences of trained length” (Cho et al., 2024b). In addition to parity, integer addition is often used as a test-bed for length generalization (McLeish et al., 2024; Cho et al., 2024b). Chang and Bisk (2025) showed that transformers show poor OOD performance for the simple task of counting (including task variants). Deletang et al. (2023) showed similar results for a series of more challenging tasks based on formal languages grouped within the Chomsky hierarchy.
@@ C.7 Length Generalization, Task Decomposition, and Scratchpads
Like the path-star task, integer arithmetic is a simple task with a simple underlying algorithm that LMs fail on. Also, a trivial reverse solution exists. Zhou et al. (2024b) shows that the reverse solution to arithmetic is more robust in terms of length generalization. Note that both these reverse solutions work because they do not need to think multiple steps ahead. For the path-star task, this betrays a lack of reasoning, however, for addition, this conforms with how humans perform arithmetic and thus feels like a more valid solution despite not requiring multi-step reasoning. The more interesting question is if models can learn to find the trivial order (which we find does not happen for the BoW experiments). See Appx. C.9 for order considerations.
@@
Yin et al. (2024) tried thinking tokens for the path-star problem to negative results. [^36] Note these are done under the original settings that allow for spurious correlations and overfitting, and thus may have failed due to other reasons. This is interesting because using $M$ thinking tokens can, in theory, provide a trivial solution by computing the reverse arm with the thinking token and then the forward arm from the reverse solution (similar to the BoW experiments, just with even less supervision). We conjecture that thinking tokens do not work for the path-star task as they do not provide additional decomposition supervision. [^37] However, these experiments need to be redone for confirmation since they did not use an online dataset.
C.8 Positional Embeddings
With length generalization comes the need to have positional embeddings that allow for exact matching across long lengths and generalize to unseen positions (Kiyono et al., 2021; Kazemnejad et al., 2023; Ruoss et al., 2023; Li et al., 2024a; McLeish et al., 2024). Chang and Bisk (2025) showed that different embedding types generalize differently to different counting tasks. Such methods will be important considerations for graph-based tasks when scaling up the size of graphs and considering length generalization. This can also lead to some unexpected results like using no positional embeddings (NoPE) being possible for decoder-only models (Irie et al., 2019; Tsai et al., 2019; Haviv et al., 2022; Chi et al., 2023; Kazemnejad et al., 2023; Wang et al., 2024b; Irie, 2024; Zuo et al., 2025) and can lead to better length generalization for symbolic reasoning tasks (Kazemnejad et al., 2023). However, Wang et al. (2024b) showed that NoPE fails to generalize due to a collapse in the attention head distribution as the context size increases. Yang et al. (2025) followed this up with a hybrid strategy that combines NoPE, for its strong token retrieval and RoPE for its inductive biases. Frydenlund (2024) found that the choice of positional embedding mattered for the path-star task and that NoPE worked when using decoder-only models.
@@ C.8 Positional Embeddings
C.9 Order Matters and Reversal Curse
Prior work has explored the impact of ordering source-side information in LLMs for premise order on the complex tasks of reasoning (Wang et al., 2023a; Chen et al., 2024c; Allen-Zhu and Li, 2024; Shah et al., 2024) and proof generation (An et al., 2024). Liu et al. (2024b) has shown that LLMs struggle to retrieve relevant information in long contexts when that information is placed in the middle of the context compared to either the front or back. Frydenlund (2024) showed that order matters for the path-star where they considered that the query should proceed the graph.
@@ C.9 Order Matters and Reversal Curse
Papadopoulos et al. (2024) discovered an asymmetry in perplexity between models trained either in the forward or reverse direction, with the forward having consistently lower perplexity. This is surprising given that both directions give a valid and theoretically equivalent factorization of the sequential probability.
@@
C.10 Non-AR, Iterative-AR, and Discrete Diffusion Models
@@
Bachmann and Nagarajan (2024) used a ‘teacher-less’ model which masks out all input tokens. Frydenlund (2024) connected this model to NAR models and also showed that the path-star task was solvable via an encoder-only model with NAR and IAR training. This was based on a modified version of the CMLM model (where the conditional ‘C’ part of the model is removed) (Ghazvininejad et al., 2019). Frydenlund (2024) incorrectly implied that the original ‘teacher-less’ model was non-causal when considering it as a NAR model. The model described by Monea et al. (2023) is meant to modify an autoregressive model post-hoc and thus is designed to keep the causal constant. However, in terms of independently modeling and predicting multiple tokens it behaves exactly like a NAR model (i.e., loss, training, and inference procedure).
@@
C.11 Future Token Prediction
Early works in future prediction designed models which could predict $N$ tokens into the future by creating $N$ separate hidden-states and training on each state with cross-entropy against a single future token (Goodman et al., 2020; Qi et al., 2020). Thus these are not truly belief-states directly, however, a belief-state must be present in the model in order to generate the $N$ separate hidden-states. The general goal of this was for improved training by explicitly learning to predict future tokens and hence plan for future tokens, and was not for multi-token inference. Heo et al. (2024) also used multi-state prediction for future n-grams but also introduced a method to explicitly create representations that are compositional into the future. Gloeckle et al. (2024) proposed an efficient training method for $N$ token prediction which repurposed $N$ multi-head attention to create the $N$ hidden-states. [^38] Qian et al. (2025) considered the use of multi-token prediction to alleviate prompt sensitivity for LLMs. Interestingly, they mentioned that the four future tokens used by Gloeckle et al. (2024) are insufficient for their task. This highlights a potential benefit of RITF, which may be able to scale to more distant future tokens since this is decoupled from any model parameterization. While they did consider how to use this to increase inference speed, their main goal was to demonstrate that training with multi-token prediction increases performance on downstream tasks even if this ability is removed during inference. Because of the success of training with multi-token prediction, it has as been adopted into foundational models such as DeepSeek3 (Liu et al., 2024a), which used sequential prediction method which requires introducing $N$ multiple parameterized modules for $N$ future tokens.
Gerontopoulos et al. (2025) also introduced a method for training a model with multi-token prediction which includes extra register tokens into the input sequence during training. These extra tokens induce creating extra hidden-states from which future tokens can be predicted. Thus, instead of modifying the model’s architecture by re-purposing heads to create extra hidden-state, this simply modifies the input sequence. Since they only use this for training, they also prevent these register tokens from being attended to by regular tokens, which would result in a inference-time bias. They showed the effectiveness of this method compared to Gloeckle et al. (2024). [^39] They also show their method works on the path-star task using a pretrained GPT2 model. Their method is similar to our RITF method, except whereas they make future predictions from extra hidden-states created by modifying the input to the model, we instead reuse hidden-states for multiple future predictions via using a structured loss function. This helps highlight the connection between inputs and targets for multi-token prediction.
Cai et al. (2024) expanded on the multi-head method for multi-token prediction by incorporating it with speculative decoding faster inference (Xia et al., 2022; Chen et al., 2023a; Leviathan et al., 2023). [^40] As an aside, speculative decoding was also the original purpose of the ’teacher-less’ model (Monea et al., 2023) and, as with NAR/IAR, the main motivating factor for speculative decoding is improved inference speed.. The parameterized modules of DeepSeek3 can also be retrained during inference and used for speculative decoding (Liu et al., 2024a).